Curriculum Vitae

Brian D. Fitzpatrick

January 11, 2012

Contact Information

Duke University Box 90320 Durham, NC 27708 USA Phone: (210)-410-7732 Fax: (919)-660-2821

                                  E-mail: bfitzpat@math.duke.edu
                                  Web: math.duke.edu/grad/bfitzpat

Education

Duke University

• Ph.D. in Mathematics, Expected May 2015

Trinity University

• B.A. in Mathematics, magna cum laude May 2010

• Advisor: Dr. Eduardo Cabral Balreira

• Thesis: On Minimal Surfaces and their Representations

Budapest Semesters in Mathematics

• Attended program in Spring 2009 and Fall 2009

Awards

Trinity University Financial Aid for Science and Technology Grant

• Received $9,200 grant in October 2009

Trinity University Mach Fellowship

• Received $2,000 fellowship in March 2009

Brian D. Fitzpatrick

Trinity University Outstanding Junior in Mathematics Award • Received in March 2009

Selected Activities

Duke University Mathematics Graduate Representative • Spring 2011

Teaching

Duke University

• Taught MATH 41L: Introduction to Calculus II With Applications, Fall 2011

• TA for MATH 25L: Labratory Calculus I, Fall 2010

Additional Coursework

Princeton University (Summer 2009)

• Introductory Topics in Fourier Analysis, Instructor: Elias Stein

• Several Complex Variables and Partial Differential Equations, Instructor: Alexander Nagel

• Partial Differential Equations in Geometry, Instructor: Gang Tian

Mathematics Competitions

Putnam Exam • Participated in 2007 (score: 1), 2008 (score: 20), and 2009 (score: 28)